Rooney Company incurs annual fixed costs of $101,420. Variable costs for Rooney’s product are $21.35 per unit, and the sales price is $35.00 per unit. Rooney desires to earn an annual profit of $46,000.


Required
Use the per unit contribution margin approach to determine the sales volume in units and dollars required to earn the desired profit. (Do not round intermediate calculations. Round your final answers to the nearest whole number.)

Sales in dollars Sales volume in units
Answer #1

Contribution margin per unit = Sales - variable cost

= $35-21.35 = $13.65

Break even point = ( Fixed cost + desired profit)/ Contribution margin per unit

= ($101420 + 46000) / 13.65 = 10800 units

Sales in dollars = 10800*35 = $378000